OpenStudy (anonymous):

Marsha wanted to explain the procedure for solving the following equation to Janice (3x+15=24) What is the explanantion that Marasha could use to explain to Janice.

OpenStudy (anonymous):

Please Help

OpenStudy (anonymous):

To solve an equation, you need solve for the variable, in this case, x. In order to do this, you first need to get 3x by itself by subtracting 15 on both sides. This gives 3x=9. Finally, in order to get x by itself, you divide by 3 on both sides, leaving x=3. Hope this helped!

OpenStudy (anonymous):

3x+15=24 3x=24-15=9 x=3

OpenStudy (anonymous):

Yes , thanks